OpenStudy (anonymous):

How would I solve the following differential equation? y'=(t-y)^2

OpenStudy (amistre64):

burnoodle it

OpenStudy (amistre64):

maybe

OpenStudy (amistre64):

y'=(t-y)^2 y'=t^2-2ty+y^2 hmm or try the wolfs method : http://www.wolframalpha.com/input/?i=y%27%3D%28t-y%29%5E2

OpenStudy (amistre64):

they appear to just sub in a new function of t: v=(t-y)

OpenStudy (anonymous):

that's pretty dope

OpenStudy (amistre64):

v' = 1-y' ; y' = -v'+1 -v'+1 = v^2

OpenStudy (amistre64):

v' = -v^2+1 \[\int (\frac{v'}{-v^2+1}=1)\] \[-\frac{1}{2}\int \frac{-2v'}{-v^2+1}dt=t+c\] i think i lose it about there